OpenStudy (a.ahmed):

prove that : (2n)! = (1x3x5x.......(2n-1)x2^n x(n)!

OpenStudy (sshayer):

\[2n!=2n(2n-1)(2n-2)(2n-3)...3.2.1=1.3.5...(2n-1).2^n[n(n-1)(n-2)...3.2.1]\] \[=1.3.5...(2n-1) .2^n.n!\]

OpenStudy (sshayer):

i write missing part again. \[2^n[n.(n-1).(n-2)...3.2.1]\] =

OpenStudy (a.ahmed):

Thank you , sshayer .and i am sorry for wasting your time

OpenStudy (sshayer):

no problem.